• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name E.Chudakov

    Log entry time 17:42:51 on November 20, 2008

    Entry number 250062

    keyword=Experiment: summary for Nov 19 (swing) - 20 (day)

    Nearly 24h lost to the reg.

    <pre>
    date ABU BANU BNAU
    2008/11/19 swing 1.42 5.33 1.23 control access 6h - Q1
    2008/11/20 owl 4.27 0.52 3.22 auxiliary measurements
    2008/11/20 day 0.00 4.31 3.69
    </pre>

    At Nov 19 about 16:00 the left Q1 started tripping more and more often.
    Jack and Todd worked on it till 22:00, then gave up.

    Nov 20 owl: a number of auxiliary measurements have been done:
    a) Moller measured -74.1+/-0.2+/-2.
    b) A number of gas Cherenkov tests were done.

    Nov 20 day: full shift in control access
    a) by the end of the shift the Q1 problem was solved by the tech's crew -
    the flow level was finally increased, and a problem
    with the power supply was identified and fixed;
    b) the dehumidifier compressor fixed;
    c) the RICH DAQ seems to be fixed, the RICH is included into the data stream;
    d) BCM, left arm, fixed;
    e) Many things fixed on the right arm;
    f) Gas Cherenkov - new test shielding installed;

    Plans:
    a) take production data;
    a) Nov 21 (day) - change the 1/2wl plate;
    b) Nov 25 (day) - target cell change.

    For the gas Cherenkov studies one may take one or two short (20min) low
    current runs per day.


    A copy of this log entry has been emailed to: transversity@jlab.org